Fermi National Laboratory


Accelerator Update - July 9-July 11

The integrated luminosity for last week was 2719.1 nb-1.

Wednesday July 9
The day shift began with Operations monitoring store 2772, with MiniBooNE taking beam, and with the Antiproton Source (Pbar) stacking.

Operations halted Booster beam at 10:50 AM, so technicians could switch in a RF station (BRF13). All Booster RF stations were back on line by 11:24 AM. Starting at the same time, a Linac expert switched from the H- source to the I- Source.

The I- chopper tripped off at 11:18 AM due to a vacuum burst. Operators reset the device.

Pbar resumed stacking at 11:26 AM.

A lightning induced power glitch hit the accelerators at 12:45 AM. The Linac beam permit went away, but the stack and store remained stable. Operations contacted an expert who had the system reset by 12:55 PM.

A Booster RF station (BRF13) tripped off at 12:58 PM. It reset okay.

Linac resumed sending beam to the other accelerators at 1:03 PM. Pbar resumed stacking.

Two Booster RF stations (BRF1 & 2) tripped off at 2:10 PM due to a vacuum beam permit.

Main Injector (MI) beam disappeared due to a RF problem at 3 PM. Beam resumed at 3:13 PM.

The Linac 400 MeV area suffered nine beam trips between 5:47 PM and 11 PM.

At 8:39 PM, a D-zero (D0) experimenter reported that one of their Roman Pots became stuck in the beam. Operation consulted with experts.

Experts began Tevatron (TeV) end-of-store studies at 10:14 PM

A MI RF station power supply tripped off for the first of five times at 10:35 PM. The trips occurred over a half hour period.

Thursday July 10
The midnight shift began with end-of-store and other accelerator studies.

A MI RF station (MIRF2) began tripping off at 12:40 AM due to conflicts with a NUMI study event. The NUMI study ended at 2 AM and the tripping stopped.

A Booster RF station (BRF8) tripped and caused an East Anode power supply trip. This happened several times so BRF8 was bypassed and left off.

Switchyard experts completed their studies at 1:10 AM.

A lightning strike at 1:21 AM caused a MI power supply to trip. It reset okay.

A Booster vacuum pump for BRF5 & 6 tripped off at 2:15 AM. Operators reset the pump, but then three other vacuum pumps tripped off. Everything was back to normal with in fifteen minutes. Beam returned to MiniBooNE and Pbar.

Booster RF station 13 & 14 tripped off due to a vacuum high pressure indication at 2:55 AM. It reset okay.

Operations terminated store 2772 at 4:09 AM.

Operators made the Collider Detector Facility (CDF) and D0 ready for access at 4:17 AM.

CHL technicians began repairing a shutoff control gas leak for the A2 wet engine.

At 5:15 AM, a TeV sector A4 LCW pump tripped off. The Duty Mechanic also reported a LCW leak at the sector A2 service building. Operators notified the water experts who fixed the A2 leak and had the A4 LCW pump back on by 6:06 AM. Unfortunately the A4 LCW pump tripped off again at 6:31 AM. Operations contacted the experts.

Power supply techs investigated the supply that got soaked by the TeV sector A2 LCW leak.

Technicians investigated the D0 Roman Pot problem and found a broken coupler at 7:39 AM. They began repairs.

Booster RF experts disconnected BRF8 at 9:20 AM from its power supply and began repairs. They finished their work at 12:45 PM.

Booster beam and stacking resumed at 9:35 AM.

Techs completed the D0 Roman Pot work at 9:38 AM.

Cryo techs reported finding a bad bearing for the TeV sector C2 wet engine. They completed their work by 11:09 AM.

Experts began pre-shot setup tuning and squeeze at 1 PM.

Operations put the TeV into shot setup at 1:35 PM and established store 2774 at 3:35 PM. Its initial luminosity was 35.36E30. Pbar resumed stacking.

Booster suffered a vacuum burst at 5:53 PM. It reset okay.

Friday July 11
The midnight shift began with Operations monitoring store 2774, with Pbar stacking, and with MiniBooNE taking beam.

A MI power supply tripped off at 12:35 AM. It tripped off twice more over the next two hours. Operators investigated but found nothing out of the ordinary. These trips interfered with stacking.

At 4:58 AM, a wet engine at TeV sector F4 slowed to near its trip point. The cryo system remained stable, but the Cryo Coordinator stated that this engine would not last long without repairs. Since turning this engine off might affect the present store, repairs were held off until later.

Today's Plan
The F4 wet engine and other maintenance work will occur after the store is terminated this morning for a planned access. Operations will put the TeV into shot setup early this afternoon. The plan for the weekend is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/11/2003   email Fermilab